OpenStudy (anonymous):

How do I determine the mole fraction of manganese(II) fluoride in a 9.25 M aqueous solution of manganese(II) fluoride. The density of the solution is 1.04 g mL-1.

OpenStudy (aaronq):

determine the moles of water and the moles of MgF2, then use the equation

OpenStudy (anonymous):

How do I find the moles.

OpenStudy (anonymous):

Do I use M=n/V with m=n*Mm

OpenStudy (aaronq):

yeah, you might also have to use \(\rho=\dfrac{m}{V}\) since you're given the density

OpenStudy (anonymous):

Thank you

OpenStudy (aaronq):

you're welcome

OpenStudy (anonymous):

Do I assume volume is = to 1L

OpenStudy (aaronq):

yeah, it makes things easier. any volume works though

OpenStudy (anonymous):

:)